Assume you work for a lending institution and one of your regular customers has approached you to borrow $200,000 to purchase a fleet of trucks for her company. Explain what it means to amortize a loan. What tools would you use to amortize a loan? Why is amortization beneficial to the lender? Why is it beneficial to the borrower?

1. Explain what it means to amortize a loan.

2. Describe the tools you would use to amortize a loan.

3. Discuss why amortization is beneficial to the lender.

4. Discuss why amortization is beneficial to the borrower.

PLEASE NO PLAGIARISM
Answer #1

1. Explain what it means to amortize a loan.

  • The regular payment is made by borrower which is called installment
  • Amortization of a loan means payment of interest on outstanding balance and proportional repayment of principal outstanding on regular intervals
  • The periodic payments or installments are fixed which includes variating payment of interest and principal repayment
  • The interest is paid on outstanding balance of loan
  • The proportionate repayment of loan is paid from regular payment after paying interest portion

Installment or Payment = Interest on Outstanding loan balance + Proportionate principal repayment which is balance amount after payment of interest

2. Describe the tools you would use to amortize a loan.

  • The following formula is used to calculate the monthly installment:

Formula: PMT = Payment = |PV| x R% x (1+R%)^N / ((1+R%)^N - 1)

I/Y = R = Rate =

FV = Future value =

N = Total payment term x Frequency =

PV =

The two different tools used ; 1) Financial calculator ; 2) Manual short hand formula as above

Assuming yearly mortgage installment; interest rate as 6%; total loan term 5 years and loan value $200,000

Using financial calculator BA II Plus - Input details:

#

I/Y = Rate/Frequency =

                         6.000000

FV = Future value =

$0.00

N = Total payment term x Frequency =

5.00

PV =

-$200,000.00

CPT > PMT = Payment =

$47,479.28

Formula: PMT = Payment = |PV| x R% x (1+R%)^N / ((1+R%)^N - 1)

PMT = 200000*0.06* (1+0.06)^5 / ((1+0.06)^5 - 1)) =

$47,479.28

Assuming repayment monthly:

Using financial calculator BA II Plus - Input details:

#

I/Y = Rate/Frequency = 6/12 =

                         0.500000

FV = Future value =

$0.00

N = Total payment term x Frequency = 5 x 12 =

60.00

PV =

-$200,000.00

CPT > PMT = Payment =

$3,866.56

Formula: PMT = Payment = |PV| x R% x (1+R%)^N / ((1+R%)^N - 1)

PMT = 200000*0.005* (1+0.005)^60 / ((1+0.005)^60 - 1)) =

$3,866.56

3. Discuss why amortization is beneficial to the lender.

  • Lender gets cash flows on regular intervals in form of installment which contains interest and proportionate principal value also
  • As and when lender gets is principal and interest value back they can further lend to other clients on different intervals on revised rates

4. Discuss why amortization is beneficial to the borrower

  • The installment helps repayment of loan value on periodic basis hence the interest on loan gets reducing with passage of time
  • Higher the frequency of installment ensures the lowest interest repayment because we end up paying loan value faster.